OpenStudy (anonymous):

Find an equation for the line satisfying the given conditions. Through (4, 4) and parallel to 5x - 4y = 4.

OpenStudy (whpalmer4):

Okay, for two lines to be parallel, they must have the same slope. Find the slope of the existing line, then use the point-slope formula to write the equation of the parallel line that goes through the known point. Your existing line is in Standard form: \[Ax + By = C, a \ne 0, b \ne 0\]Rearrange it into slope-intercept form and read off the value of the slope (\(m\)): \[y = mx + b\] The point-slope formula for a line through a known point \((x_0, y_0)\) is:\[y-y_0 = m(x-x_0)\]

OpenStudy (anonymous):

Thank you!

OpenStudy (whpalmer4):

What did you get for your answer?